# Manifold-aware Representation Learning for Degradation-agnostic Image Restoration

### v1

@@ B.4 Ablation Regarding the Optimization Objectives
Tab. C shows that replacing SPD-based contrastive learning with a standard Euclidean-space contrastive loss (w/o SPD) results in a clear performance drop, demonstrating the advantage of modeling second-order channel correlations on the SPD manifold rather than relying solely on first-order vector similarities. When the entire contrastive module is removed (w/o CL & SPD), performance degrades even further, indicating that aligning shallow and deep features is essential for effective representation learning. Moreover, removing the Fourier loss (w/o Fourier Loss) slightly reduces performance, suggesting that frequency-domain supervision provides additional benefits. Overall, the full model achieves the best results, confirming the effectiveness of jointly optimizing spatial, frequency, and SPD-manifold-based structural consistency. Note that throughout all the experiments, we set $\lambda_{ctrs}$ = 0.05 and $\lambda_{ctrs}$=0.1.
B.5 Shallow-Latent Feature Similarity
@@
Appendix C Additional Visual Results.
C.1 3 Degradation
@@
For the 5-degradation setting, we provide visual comparisons for the low-light enhancement task in Fig. D. As illustrated, the proposed MIRAGE produces noticeably cleaner outputs with improved luminance restoration and better color consistency compared to MoCE-IRzamfir2025moce , demonstrating its robustness under challenging illumination conditions.
C.3 Composited Degradation
@@
Fig. E and Fig. F present visual comparisons under more challenging composite degradations, namely low-light + haze + snow and low-light + haze + rain, respectively. As observed, our method reconstructs significantly more scene details and preserves structural consistency, whereas MoCE-IR zamfir2025moce tends to produce noticeable artifacts and over-smoothed regions under these complex conditions.
C.4 Zero-Shot Underwater Image Enhancement
@@
Appendix D Limitations and Future Work
While the proposed MIRAGE achieves new state-of-the-art performance on most all-in-one image restoration benchmarks, we observe that its deblurring performance still lags slightly behind MoCE-IR zamfir2025moce . We attribute this to the relatively compact model size of our current design, which favors efficiency over aggressive capacity. To address this, future work will explore scaling up the model size to be on par with larger architectures such as PromptIR potlapalli2023promptir , MoCE-IR zamfir2025moce , and AdaIR cui2025adair , aiming to further boost performance while maintaining the architectural elegance and efficiency of our design. Moreover, our current SPD-based contrastive learning leverages a conventional InfoNCE loss in Euclidean space after projecting SPD features. While effective, it does not fully exploit the intrinsic geometry of the SPD manifold. As part of future efforts, we plan to investigate geodesic-based contrastive formulations and Riemannian-aware optimization strategies, which may offer a more principled and theoretically grounded way to align structured representations across semantic scales.
Appendix E Broader Impact
Image restoration (IR) is a foundational task with widespread applications across photography, remote sensing, surveillance, autonomous driving, medical imaging, and scientific visualization. By proposing a unified and efficient framework capable of handling diverse degradation types with minimal computational cost, our work has the potential to benefit a wide range of real-world scenarios where image quality is compromised due to environmental or hardware constraints. The lightweight design of MIRAGE enables deployment on resource-limited edge devices, such as mobile phones, drones, or embedded cameras in IoT systems. This democratizes access to high-quality image restoration, which may positively impact users in low-resource settings or in critical applications like emergency response and environmental monitoring. From a research perspective, our modular design and SPD-based contrastive formulation may inspire further work on principled representation learning for restoration and beyond, encouraging more geometrically-aware approaches in low-level vision. While our method is intended for general-purpose image enhancement, we acknowledge that improved image restoration techniques can also be misused for deceptive media editing or surveillance applications that raise privacy concerns. We encourage future practitioners to adopt this technology responsibly and in alignment with ethical standards. Our code and models will be released with appropriate licenses and documentation to promote transparency and responsible use.

### v2

@@ C.4 Ablation Regarding the Optimization Objectives
demonstrating the advantage of modeling second-order channel correlations on the SPD manifold rather than relying solely on first-order vector similarities. When the entire contrastive module is removed (w/o CL & SPD), performance degrades even further, indicating that aligning shallow and deep features is essential for effective representation learning. Moreover, removing the Fourier loss (w/o Fourier Loss) slightly reduces performance, suggesting that frequency-domain supervision provides additional benefits. Overall, the full model achieves the best results, confirming the effectiveness of jointly optimizing spatial, frequency, and SPD-manifold-based structural consistency. Note that throughout all the experiments, we set $\lambda_{ctrs}$ = 0.05 and $\lambda_{ctrs}$=0.1.
C.5 Shallow-Latent Feature Similarity
@@
C.6 More Generlization Evaluation
Table D: Zero-shot evaluation on real-world under-display camera datasets TOLED and POLED (Zhou et al., 2021).
Method TOLED (PSNR / SSIM / LPIPS) POLED (PSNR / SSIM / LPIPS)
AirNet (Li et al., 2022) 14.58 / 0.609 / 0.445 7.53 / 0.350 / 0.820
PromptIR (Potlapalli et al., 2024) 16.70 / 0.688 / 0.422 13.16 / 0.583 / 0.619
DiffUIR (Zheng et al., 2024a) 29.55 / 0.887 / 0.281 15.62 / 0.424 / 0.505
MIRAGE-S (Ours) 28.01 / 0.881 / 0.293 16.93 / 0.604 / 0.500
To further assess generalization beyond synthetic settings, we evaluate MIRAGE-S on the real-world TOLED and POLED under-display camera datasets (Zhou et al., 2021). As shown in Tab. D, MIRAGE-S achieves strong performance across both benchmarks. On POLED, which contains more severe signal attenuation and non-linear spatial artifacts, MIRAGE-S clearly surpasses prior methods across all three metrics, indicating robust transfer to challenging real-world degradations. On TOLED, MIRAGE-S remains competitive and delivers results close to diffusion-based DiffUIR despite its significantly lower complexity. These findings suggest that the proposed mixed-backbone architecture and SPD-based alignment maintain good stability under real sensor degradations and generalize reliably across distinct UDC hardware conditions.
Appendix D Additional Visual Results.
D.1 3 Degradation
@@ C.6 More Generlization Evaluation
For the 5-degradation setting, we provide visual comparisons for the low-light enhancement task in Fig. D. As illustrated, the proposed MIRAGE produces noticeably cleaner outputs with improved luminance restoration and better color consistency compared to MoCE-IRZamfir et al. (2025), demonstrating its robustness under challenging illumination conditions.
D.3 Composited Degradation
@@
Fig. E and Fig. F present visual comparisons under more challenging composite degradations, namely low-light + haze + snow and low-light + haze + rain, respectively. As observed, our method reconstructs significantly more scene details and preserves structural consistency, whereas MoCE-IR Zamfir et al. (2025) tends to produce noticeable artifacts and over-smoothed regions under these complex conditions.
D.4 Zero-Shot Underwater Image Enhancement
@@
Appendix E Limitations and Future Work
While the proposed MIRAGE achieves new state-of-the-art performance on most all-in-one image restoration benchmarks, we observe that its deblurring performance still lags slightly behind MoCE-IR Zamfir et al. (2025). We attribute this to the relatively compact model size of our current design, which favors efficiency over aggressive capacity. To address this, future work will explore scaling up the model size to be on par with larger architectures such as PromptIR Potlapalli et al. (2024), MoCE-IR Zamfir et al. (2025), and AdaIR Cui et al. (2025), aiming to further boost performance while maintaining the architectural elegance and efficiency of our design. Moreover, our current SPD-based contrastive learning leverages a conventional InfoNCE loss in Euclidean space after projecting SPD features. While effective, it does not fully exploit the intrinsic geometry of the SPD manifold. As part of future efforts, we plan to investigate geodesic-based contrastive formulations and Riemannian-aware optimization strategies, which may offer a more principled and theoretically grounded way to align structured representations across semantic scales. Additionally, different degradations may favor different proportions of convolution, attention, and MLP capacity. Learning such ratios dynamically is an interesting direction and could further adapt MIRAGE to degradation-specific characteristics. We view this as a promising avenue for future research.
Appendix F Broader Impact
Image restoration (IR) is a fundamental task with applications in photography, remote sensing, surveillance, autonomous driving, medical imaging, and scientific visualization. By proposing a unified and efficient framework capable of handling diverse degradation types with minimal computational cost, our work may benefit scenarios where image quality is compromised by environmental or hardware constraints. The lightweight design of MIRAGE further enables deployment on resource-limited devices such as mobile phones, drones, or embedded cameras, which can support use cases in low-resource settings or critical domains like emergency response and environmental monitoring. From a research perspective, our modular design and SPD-based contrastive formulation may encourage further exploration of geometrically-aware representation learning in restoration and related areas.
Appendix G Use of Large Language Models (LLMs)
We used OpenAI’s GPT-based Large Language Models (LLMs) (OpenAI, 2023; 2022) to polish the writing and improve the readability of the paper. The models were not used for developing the methodology, running experiments, or analyzing results. All scientific contributions remain entirely the work of the authors.